Find the constant term that should be added to make the following a perfect square trinomial:
x^2 + 7x
Multiple choice:
a)49
b)49/2
c)7/2
d)49/4
Thank you for your help

Found 2 solutions by stanbon, Nate:
Answer by stanbon(75887)

Find the constant term that should be added to make the following a perfect square trinomial:
x^2 + 7x
Square (1/2) the coefficient of the x term.
(7/2)^2 = 49/4
----------
Cheers,
Stan H.

Answer by Nate(3500)

(x + k)^2 = x^2 + 2k*x + k^2
2k = 7
k = 3.5
k^2 = 3.5^2 = 12.25 = 49/4
x^2 + 7x + 49/4